The Perimeter of a rectangular garden is 24ft. The length is 2 more feet than the width. Find the length and the width of the garden.
Answer by checkley71(8403)

L=W+2
P=2L+2W
24=2(W+2)+2W
24=2W+4+2W
4W=24-4
4W=20
W=20/4
W=5 FT. IS THE WIDTH OF THE GARDEN.
L=5+2)
L=7
PROOF:
2+7+2*5=14=10=24